OpenStudy (anonymous):

I'm not sure where to start... Please help. assume that the sales of a certain appliance dealer are approximated by a linear function. Suppose that sales were $5000 in 1982 and $73500 in 1987. Let x+0 represent 1982. Find the equation giving yearly sales S(x)

OpenStudy (anonymous):

point is (0,5000),(5,73500) y-5000=13700(x-0) y=13700x+5000

OpenStudy (anonymous):

thank you